# FGAIF: Aligning Large Vision-Language Models with Fine-grained AI Feedback

### v1

@@ 1 Introduction
Large Language Models (LLMs) like GPT-3 (Brown et al., 2020) and ChatGPT (OpenAI, 2022) have showcased remarkable abilities in language processing. However, their ability to handle multimodal inputs combining both visual and textual data remains inadequate. This limitation has drawn research attention to Large Vision-Language Models (LVLMs) which achieve massive success in various vision and language tasks (e.g. Visual Question Answering (Antol et al., 2015) and Image Captioning (Lin et al., 2014)).
Although LVLMs have achieved significant success in tasks requiring visual-textual understandings, the challenge of misalignment between vision and language modalities Sun et al. (2023) has not been solved, leading to “hallucination” in generated textual responses (Jing et al., 2023). As shown in Figure 1, there are three kinds of hallucinations in the context of LVLMs, including (1) Object Existence Hallucination, where non-existent objects are mistakenly referenced; (2) Object Attribute Hallucination, involving inaccuracies in the depiction of object attributes like color, shape, and size; and (3) Object Relationship Hallucination, where the descriptions inaccurately portray the interactions or spatial relationships between objects, leading to misrepresentations of their positions, interactions, and actions involving two or more objects Jing et al. (2023); Zhai et al. (2023). Therefore, mitigating the hallucinations and generating faithful responses are key to building practical applications of LVLMs.
@@
Our contribution can be summarized as follows: 1) We propose a novel fine-grained artificial intelligence-based hallucination labeling method, which can detect three types of hallucinations (i.e., object existence, object attribute, and object relation) in terms of sub-sentence level and eliminate the need for manual annotation. 2) To the best of our knowledge, we are the first to provide multiple types and segment-level feedback towards modalities alignment in LVLMs, which can mitigate three kinds of hallucination in LVLMs. 3) We conduct comprehensive experiments on several hallucination benchmarks and one general benchmark. The experimental results demonstrate the effectiveness of FGAIF. In addition, the ablation study shows the necessity of each module in FGAIF. As a byproduct, we have released our code and parameters111https://github.com/du-nlp-lab/FGAIF..
@@ 2 Related Work
Large Vision-Language Model
@@
Hallucinations in LVLMs
Motivated the hallucination in LLMs, more researchers shifted research attention to hallucination in LVLMs. Hallucination in the context of LVLMs is the inconsistent content between the generated response and the input image. To evaluate the hallucination in LVLMs, some work devised metrics to measure the hallucination in the response, such as FaithScore (Jing et al., 2023), CHAIR (Rohrbach et al., 2018), POPE (Li et al., 2023d), and NOPE (Lovenia et al., 2023). Recently, there have been works to mitigate hallucinations in LVLMs utilizing various technologies, such as decoding approaches (Leng et al., 2023; Huang et al., 2023), post-processing (Zhou et al., 2023; Yin et al., 2023), and construction of the higher-quality dataset (Liu et al., 2023a; Li et al., 2023c). To address the challenge of aligning image and text modalities within LVLMs and to mitigate the issue of hallucination, existing strategies offer partial solutions but lack direct guidance for modality alignment. Therefore, some research efforts (Li et al., 2023b; Yu et al., 2023; Zhou et al., 2024) have embraced the use of reinforcement learning for direct modality alignment. For example, Sun et al. (2023) developed the LLaVA-RLHF model, harnessing human-annotated preference data to minimize hallucinations in LLaVA.
Motivated by the fine-grained RL (Wu et al., 2023) and AI-based RL (Lee et al., 2023; Bai et al., 2022) methods, we propose to align modalities in LVLMs with fine-grained AI feedback. Different from existing work which needs human annotation and only provides coarse-grained feedback, our method provides fine-grained rewards and learns from AI automatic feedback.
@@ 3 Problem Formulation
Finally, we utilize well-trained reward models and a set of images and the corresponding prompts . to fine-tune the LVLM to make it generate faithful responses as , where is final optimized parameters of the LVLM . We also omit the index in this equation. is the size of data for finetuning LVLMs.
@@ 4.1 AI-based Feedback Collection
To achieve this, we first sample responses from the backbone LVLM as depicted in Section 3. Inspired by the existing fine-grained evaluation work (Jing et al., 2023; Min et al., 2023), we devise a fine-grained AI-based feedback collection method. In particular, we utilize AI models to annotate three kinds of hallucinations (i.e., object existence hallucination, object attribute hallucination, and object relationship hallucination) on the sub-sentence level for the response. In particular, to get the hallucination labels for each sub-sentence, we first split the response from the LVLM into sub-sentences as follows,
where is the -th sub-sentence of the response. Thereafter, to accurately annotate three kinds of hallucination in the sub-sentence, we extract three kinds of atomic facts (Jing et al., 2023): object existence, object attribute, and object relationship atomic facts, from the sub-sentence, using ChatGPT as follows,
where , and denote the -th object existence, object attribute, and object relation types of atomic fact derived from the sub-sentence, respectively. And is the total number of object existence/attribute/relation atomic facts for the sub-sentence. Here we omit the index of the sub-sentence for simplicity. Atomic fact is the minimal information unit and we show some examples in Appendix A. is a prompt that can instruct ChatGPT to generate three kinds of atomic facts, and corresponding details can be found in Appendix A.
@@
where , and denote the hallucination label of -th atomic fact in the sub-sentence in terms of object existence, object attribute, and object relationship types of atomic facts, respectively. is set to 1 when the output of LLaVA 1.5 indicates that the input image and the atomic fact are inconsistent (i.e., the corresponding atomic fact is a hallucination), otherwise, it is set to 0. is the prompt that can be used to prompt the LLaVA 1.5 to annotate hallucination and it is shown in Applendix A.
@@
The reason why we use LVLM to verify the consistency between atomic fact and image even if the LVLM may also introduce hallucination: Our method converts the AI labeling task into a discriminative task that usually generates a short response, and this kind of task tends not to generate hallucination, which has been demonstrated in existing work (Jing et al., 2023; Min et al., 2023). Therefore, our AI-based feedback collection method can reduce the hallucination as much as possible.
@@ 4.3 Reinforcement Learning with Fine-grained Reward
Fine-tuning language models with reinforcement learning is an effective approach to align modalities in LVLMs. To make LVLMs generate more faithful responses rather than hallucinated responses, we also resort to reinforcement learning to further fine-tune LVLMs with the fine-grained reward (right of Figure 2). Specifically, we first segment the generated response from the LVLM into sub-sentences . Then we get all kinds of rewards for each sub-sentence based on the well-trained reward model by cross-entropy loss. We define , , and as the object existence, object attribute, and object relation rewards for the -th sub-sentence. Then we have a combined reward function for each token as , where is the timestep for the last token of . is the indicator function. is a weight assigned to rewards. Thereafter, we utilize the PPO algorithm to train the policy model (i.e., the LVLM) following the existing work (Sun et al., 2023).
@@ 5 Experiment
5.1 Experimental Details
To ensure a fair and equitable comparison, we utilized same base model with the LLaVA-RLHF model whose network architecture is LLaVA. In addition, we also adopt the same architecture (i.e., LLaVA) with LLaVA-RLHF for the reward model. We compared our method with these models that used the same model backbone as ours (i.e., LLaVA (Liu et al., 2023c) and LLaVA-RLHF). We also introduced some methods with the same backbone architecture but a larger model size (i.e., LLaVA and LLaVA-RLHF). Besides, we further incorporated more advanced LVLMs for comparison, i.e., MiniGPT-4 (Zhu et al., 2023), mPLUG-Owl (Ye et al., 2023), InstructBLIP (Dai et al., 2023), and InstructBLIP.
To verify the effectiveness of our proposed FGAIF, we compare our method with baselines on several benchmarks, including QA-based hallucination benchmarks POPE (Li et al., 2023d) and MMHal-Bench (Sun et al., 2023), hallucination metrics CHAIR (Rohrbach et al., 2018) and FaithScore (Jing et al., 2023), and the general benchmark LLaVA-Bench (Liu et al., 2023c). More detailed setups for dataset and model training are shown in Appendix B.
@@ 5.1 Experimental Details
| MiniGPT-4 | 79.7 | 80.2 | 69.7 | 73.0 | 65.2 | 70.4 | 74.5 | 60.8 |
| mPLUG-Owl | 54.0 | 68.4 | 50.9 | 66.9 | 50.7 | 66.8 | 67.2 | 97.6 |
| InstructBLIP | 88.6 | 89.3 | 79.7 | 80.2 | 65.2 | 70.4 | 80.0 | 59.0 |
| InstructBLIP | 88.7 | 89.3 | 81.4 | 83.5 | 74.4 | 78.5 | 83.7 | 62.2 |
| LLaVA | 50.4 | 66.6 | 49.9 | 66.4 | 49.7 | 66.3 | 66.4 | 99.2 |
| LLaVA | 73.7 | 78.8 | 73.6 | 78.2 | 67.2 | 74.4 | 77.1 | 73.7 |
| LLaVA-RLHF | 84.8 | 83.3 | 83.3 | 81.8 | 80.7 | 79.5 | 81.5 | 41.8 |
| LLaVA-RLHF | 85.2 | 83.5 | 83.9 | 81.8 | 82.3 | 80.5 | 81.9 | 39.0 |
| FGAIF | 87.0 | 86.7 | 84.0 | 83.7 | 79.6 | 79.9 | 83.4 | 48.3 |
@@
| MiniGPT-4 | 3.39 | 0.24 | 3.0 | 2.54 | 3.67 | 80.5 | 74.5 | 81.6 | 78.9 |
| mPLUG-Owl | 2.49 | 0.43 | 0.33 | 2.58 | 1.5 | 78.7 | 46.0 | 47.4 | 57.5 |
| InstructBLIP | 2.10 | 0.58 | 2.08 | 2.67 | 2.17 | 95.4 | 96.3 | 99.1 | 97.0 |
| InstructBLIP | 2.14 | 0.58 | 1.75 | 2.82 | 2.5 | 90.9 | 91.7 | 109.3 | 97.2 |
| LLaVA | 1.55 | 0.76 | 0.00 | 1.25 | 2.00 | 75.1 | 75.4 | 92.3 | 81.0 |
| LLaVA | 1.11 | 0.84 | 0.00 | 1.13 | 1.5 | 87.2 | 74.3 | 92.9 | 84.9 |
| LLaVA-RLHF | 2.04 | 0.68 | 1.83 | 2.42 | 2.25 | 93.0 | 79.0 | 109.5 | 94.1 |
| LLaVA-RLHF | 2.53 | 0.57 | 2.67 | 2.79 | 2.33 | 93.9 | 82.5 | 110.1 | 95.6 |
| FGAIF | 3.09 | 0.36 | 3.58 | 3.21 | 3.33 | 98.2 | 93.6 | 110.0 | 100.1 |
5.2 On Model Comparison (RQ1)
The results on QA-based hallucination benchmarks (i.e., POPE and MMHal-Bench) are summarized in Table 1 and Table 2. From this table, we have several observations. (1) LLaVA and InstructBLIP performs worse than LLaVA and InstructBLIP on most cases, respectively. Compared with LLaVA, LLaVA has a strong hallucination problem, especially its over-confident problem on POPE. This indicates the importance of model size. (2) LLaVA-RLHF is better than LLaVA, which indicates the superiority of further fine-tuning with human feedback. Notably, LLaVA-RLHF even has a better performance compared to LLaVA, even though the latter has specifically more parameters. (3) Our model consistently performs better than the previous advanced in terms of all metrics and testing sets. This verifies that fine-grained artificial intelligence feedback also can be beneficial for hallucination mitigation in LVLMs. (4) Our FGAIF surpasses LLaVA-RLHF across all metrics. This implies the advantage of fine-grained artificial intelligence feedback compared to human feedback. (5) To further understand the performance of our FGAIF, we split the MMHal-Bench into three classes based on the original dataset: a) object existence (class “adversarial object”), b) object attribute (classes “object attribute” and “counting”), and c) object relation (class “spatial relation”). We observe that our method consistently achieves the best performance across all question categories.
@@ 5.2 On Model Comparison (RQ1)
| CHAIR | CHAIR | F-Score | F-Score | ||
| MiniGPT-4 | 9.4 | 17.4 | 63.9 | 61.8 | 245.1 |
| mPLUG-Owl | 6.2 | 9.5 | 85.6 | 65.7 | 75.2 |
| InstructBLIP | 2.4 | 3.8 | 93.6 | 80.0 | 45.6 |
| InstructBLIP | 2.7 | 4.0 | 94.1 | 80.8 | 46.3 |
| LLaVA | 9.1 | 22.0 | 88.9 | 72.3 | 216.0 |
| LLaVA | 10.3 | 19.8 | 87.9 | 68.3 | 121.0 |
| LLaVA-RLHF | 4.6 | 7.0 | 89.3 | 71.1 | 58.8 |
| LLaVA-RLHF | 7.7 | 20.3 | 89.7 | 73.8 | 413.8 |
| FGAIF | 3.9 | 6.2 | 91.2 | 74.7 | 60.2 |
We further show the performance of our FGAIF and baselines on hallucination metrics CHAIR and FaithScore in Table 3. InstructBLIP and InstructBLIP achieve the best performance in CHAIR and FaithScore metrics. The potential reason is that these two models tend to generate short answers and these two metrics just measure the precision of faithfulness but do not contain recall of faithfulness. Despite this, our FGAIF still outperforms the RLHF-based baseline (i.e., LLaVA-RLHF) whose answers are shorter than FGAIF, which verifies the superiority of our method.
In addition, Table 2 shows the comprehensive performance comparison of our FGAIF and the baseline methods on the general benchmark LLaVA-Bench. From this table, we observed that most models perform worst on the “Detail” (i.e., detailed description) subset and perform best on the “Complex” (i.e., complex questions) subset. This may be due to the reason that the “Detail” (i.e., detailed description) subset has more stringent requirements for faithfulness because all the content of the response is required to be an accurate description of the input image. On the contrary, the “Complex” (i.e., complex questions) subset often explores the extended content of an image, sometimes leading to open-ended discussions. Therefore, the demand for strict consistency with the image isn’t as critical. In addition, we found that the RLHF can boost the LVLM’s performance on the whole LLaVA-Bench from 81.0 (LLaVA) to 94.1 (LLaVA-RLHF). Furthermore, our FGAIF can bring more performance gain in terms of the “Conv” subset, “Detail”, “Complex” subset, and full set), compared with LLaVA-RLHF. This further indicates the advance of our method.
5.3 On Ablation Study (RQ2)
To verify the effect of each component in our FGAIF, we devise the following variant methods for ablation study: 1) w/o-Obj: To demonstrate the effect of the object hallucination feedback, we remove the object existence reward model in this method; 2) w/o-Att: To show the necessity of the attribute hallucination feedback, we remove the object attribute reward model in this method; 3) w/o-Rel: To demonstrate the effect of the relation hallucination feedback, we remove the object relation reward model in this method; 4) w/o-AIF: To show the benefit of using reinforcement learning from fine-grained artificial intelligence feedback, we remove all the reinforcement learning components in this variant; 5) w-Coarse: To verify the advance of the fine-grained feedback compared with the traditional coarse-grained uni reward model, we replace the three fine-grained reward models with one reward model which also is trained with AI annotated data and the training phrase is the same as the previous work (Sun et al., 2023).
@@ 5.3 On Ablation Study (RQ2)
| CHAIR | CHAIR | F-Score | F-Score | F1 | Over | Hal | |
| FGAIF | 3.9 | 6.2 | 91.2 | 74.7 | 83.4 | 3.09 | 0.36 |

### v2

@@ 1 Introduction
Large Language Models (LLMs) like GPT-3 (Brown et al., 2020) and ChatGPT (OpenAI, 2022) have showcased remarkable abilities in language processing. However, their ability to handle multimodal inputs combining both visual and textual data remains inadequate. This limitation has drawn research attention to Large Vision-Language Models (LVLMs) which achieve massive success in various vision and language tasks (e.g. Visual Question Answering (Antol et al., 2015), Visual Entailment (Zhang et al., 2025), and Image Captioning (Lin et al., 2014)).
Although LVLMs have achieved significant success in tasks requiring visual-textual understandings, the challenge of misalignment between vision and language modalities (Sun et al., 2023) has not been solved, leading to “hallucination” in generated textual responses (Jing et al., 2024). As shown in Figure 1, there are three kinds of hallucinations in the context of LVLMs, including (1) Object Existence Hallucination, where non-existent objects are mistakenly referenced; (2) Object Attribute Hallucination, involving inaccuracies in the depiction of object attributes like color, shape, and size; and (3) Object Relationship Hallucination, where the descriptions inaccurately portray the interactions or spatial relationships between objects, leading to misrepresentations of their positions, interactions, and actions involving two or more objects Jing et al. (2024); Zhai et al. (2023); Zhang et al. (2024). Therefore, mitigating the hallucinations and generating faithful responses are key to building practical applications of LVLMs.
@@
Our contribution can be summarized as follows:
We propose a novel fine-grained artificial intelligence-based hallucination labeling method, which can detect three types of hallucinations (i.e., object existence, object attribute, and object relation) in terms of sub-sentence level and eliminate the need for manual annotation.
To the best of our knowledge, we are the first to provide multiple types and segment-level feedback towards modalities alignment in LVLMs, which can mitigate three kinds of hallucination in LVLMs.
We conduct comprehensive experiments on several hallucination benchmarks and one general benchmark111We released our code via https://github.com/LiqiangJing/FGAIF.. The experimental results demonstrate the effectiveness of FGAIF. In addition, the ablation study shows the necessity of each module in FGAIF.
@@ 2 Related Work
2.1 Large Vision-Language Model
@@
2.2 Hallucinations in LVLMs
Motivated the hallucination in LLMs, more researchers shifted research attention to hallucination in LVLMs. Hallucination in the context of LVLMs is the inconsistent content between the generated response and the input image. To evaluate the hallucination in LVLMs, some work devised metrics to measure the hallucination in the response, such as FaithScore (Jing et al., 2024), CHAIR (Rohrbach et al., 2018), POPE (Li et al., 2023d), and NOPE (Lovenia et al., 2023). Recently, there have been works to mitigate hallucinations in LVLMs utilizing various technologies, such as decoding approaches (Leng et al., 2023; Huang et al., 2023), post-processing (Zhou et al., 2023a; Yin et al., 2023; Chang et al., 2024), and construction of the higher-quality dataset (Liu et al., 2023a; Li et al., 2023c). To address the challenge of aligning image and text modalities within LVLMs and to mitigate the issue of hallucination, existing strategies offer partial solutions but lack direct guidance for modality alignment. Therefore, some research efforts (Li et al., 2023b; Yu et al., 2023; Zhou et al., 2024) have embraced the use of reinforcement learning for direct modality alignment. For example, Sun et al. (2023) developed the LLaVA-RLHF model, harnessing human-annotated preference data to minimize hallucinations in LLaVA.
Motivated by the fine-grained RL (Wu et al., 2023; Ramé et al., 2023; Jang et al., 2023; Zhou et al., 2023b; Wang et al., 2024) and AI-based RL (Lee et al., 2023; Bai et al., 2022) methods, we propose to align modalities in LVLMs with fine-grained AI feedback. Different from existing work which needs human annotation and only provides coarse-grained feedback, our method provides fine-grained rewards and learns from AI automatic feedback.
@@ 3 Problem Formulation
Finally, we utilize well-trained reward models and the images and the corresponding prompt . to fine-tune the LVLM to make it generate faithful responses as , where is final optimized parameters of the LVLM . We also omit the index in this equation.
@@ 4.1 AI-based Feedback Collection
To achieve this, we first sample responses from the backbone LVLM as depicted in Section 3. Inspired by the existing fine-grained evaluation work (Jing et al., 2024; Min et al., 2023), we devise a fine-grained AI-based feedback collection method. In particular, we utilize AI models to annotate three kinds of hallucinations (i.e., object existence hallucination, object attribute hallucination, and object relationship hallucination) on the sub-sentence level for the response. In particular, to get the hallucination labels for each sub-sentence, we first split the response from the LVLM into sub-sentences as follows,
where is the -th sub-sentence of the response. Thereafter, to accurately annotate three kinds of hallucination in the sub-sentence, we extract three kinds of atomic facts (Jing et al., 2024): object existence, object attribute, and object relationship atomic facts, from the sub-sentence, using ChatGPT as follows,
where , and denote the -th object existence, object attribute, and object relation types of atomic fact derived from the sub-sentence, respectively. is a prompt that can instruct ChatGPT to generate three kinds of atomic facts, as shown in Figure 3.
@@
where , and denote the hallucination label of -th atomic fact in the sub-sentence in terms of object existence, object attribute, and object relationship types of atomic facts, respectively. is set to 1 when the output of LLaVA 1.5 indicates that the input image and the atomic fact are inconsistent (i.e., the corresponding atomic fact is a hallucination), otherwise, it is set to 0. is the prompt that can be used to prompt the LLaVA 1.5 to annotate hallucination and it is shown in Figure 4.
@@
The reason why we use LVLM to verify the consistency between atomic fact and image even if the LVLM may also introduce hallucination: Our method converts the AI labeling task into a discriminative task that usually generates a short response, and this kind of task tends not to generate hallucination, which has been demonstrated in existing work (Jing et al., 2024; Min et al., 2023). Therefore, our AI-based feedback collection method can reduce the hallucination as much as possible.
@@ 4.3 Reinforcement Learning with Fine-grained Reward
Fine-tuning language models with reinforcement learning is an effective approach to align modalities in LVLMs. To make LVLMs generate more faithful responses rather than hallucinated responses, we also resort to reinforcement learning to further fine-tune LVLMs with the fine-grained reward (right of Figure 2). Specifically, we first segment the generated response from the LVLM into sub-sentences . Then we get all kinds of rewards for each sub-sentence based on the well-trained reward model by cross-entropy loss. We define , , and as the object existence, object attribute, and object relation rewards for the -th sub-sentence. Then we have a combined reward function for each token as , where is the timestep for the last token of . is the indicator function. is a weight assigned to rewards. Thereafter, we utilize the PPO algorithm Schulman et al. (2017) to train the policy model (i.e., the LVLM) following the existing work (Sun et al., 2023).
PPO is an actor-critic RL algorithm that is widely used in previous RLHF work to optimize the policy model against a reward model of human feedback. It uses a value model to estimate the value of state , and optimizes the policy model with a PPO clipped surrogate training objective. The advantage at timestep is estimated by a generalized advantage estimation function (Schulman et al., 2016): , with as a hyperparameter and as the discounting factor for rewards. is the reward assigned to , which in our case is acquired using one or multiple learned reward models. The value model is optimized with an expected squared-error loss with the value target as , where is the lagging value model. Finally, PPO is trained to optimize both policy () and value () models with their respective objectives. No reward model is being optimized during PPO training. See Appendix C for more details.
@@ 5 Experiment
MiniGPT-47B | 79.7 | 80.2 | 69.7 | 73.0 | 65.2 | 70.4 | 74.5 | 60.8 |
mPLUG-Owl7B | 54.0 | 68.4 | 50.9 | 66.9 | 50.7 | 66.8 | 67.2 | 97.6 |
InstructBLIP7B | 88.6 | 89.3 | 79.7 | 80.2 | 65.2 | 70.4 | 80.0 | 59.0 |
InstructBLIP13B | 88.7 | 89.3 | 81.4 | 83.5 | 74.4 | 78.5 | 83.7 | 62.2 |
LLaVA7B | 50.4 | 66.6 | 49.9 | 66.4 | 49.7 | 66.3 | 66.4 | 99.2 |
LLaVA13B | 73.7 | 78.8 | 73.6 | 78.2 | 67.2 | 74.4 | 77.1 | 73.7 |
LLaVA-RLHF7B | 84.8 | 83.3 | 83.3 | 81.8 | 80.7 | 79.5 | 81.5 | 41.8 |
LLaVA-RLHF13B | 85.2 | 83.5 | 83.9 | 81.8 | 82.3 | 80.5 | 81.9 | 39.0 |
FGAIF7B | 87.0 | 86.7 | 84.0 | 83.7 | 79.6 | 79.9 | 83.4 | 48.3 |
@@
MiniGPT-47B | 3.39 | 0.24 | 3.0 | 2.54 | 3.67 | 80.5 | 74.5 | 81.6 | 78.9 |
mPLUG-Owl7B | 2.49 | 0.43 | 0.33 | 2.58 | 1.5 | 78.7 | 46.0 | 47.4 | 57.5 |
InstructBLIP7B | 2.10 | 0.58 | 2.08 | 2.67 | 2.17 | 95.4 | 96.3 | 99.1 | 97.0 |
InstructBLIP13B | 2.14 | 0.58 | 1.75 | 2.82 | 2.5 | 90.9 | 91.7 | 109.3 | 97.2 |
LLaVA7B | 1.55 | 0.76 | 0.00 | 1.25 | 2.00 | 75.1 | 75.4 | 92.3 | 81.0 |
LLaVA13B | 1.11 | 0.84 | 0.00 | 1.13 | 1.5 | 87.2 | 74.3 | 92.9 | 84.9 |
LLaVA-RLHF7B | 2.04 | 0.68 | 1.83 | 2.42 | 2.25 | 93.0 | 79.0 | 109.5 | 94.1 |
LLaVA-RLHF13B | 2.53 | 0.57 | 2.67 | 2.79 | 2.33 | 93.9 | 82.5 | 110.1 | 95.6 |
FGAIF7B | 3.09 | 0.36 | 3.58 | 3.21 | 3.33 | 98.2 | 93.6 | 110.0 | 100.1 |
5.1 Experimental Details
To ensure a fair and equitable comparison, we utilized same base model with the LLaVA-RLHF model whose network architecture is LLaVA7B and underwent a supervised fine-tuning stage. In addition, we also adopt the same architecture (i.e., LLaVA13B) with LLaVA-RLHF for the reward model. We compared our method with these models that used the same model backbone as ours (i.e., LLaVA7B (Liu et al., 2023c) and LLaVA-RLHF7B). We also introduced some methods with the same backbone architecture but a larger model size (i.e., LLaVA13B and LLaVA-RLHF13B). Besides, we further incorporated more advanced LVLMs for comparison, i.e., MiniGPT-47B (Zhu et al., 2023), mPLUG-Owl7B (Ye et al., 2023), InstructBLIP7B (Dai et al., 2023), and InstructBLIP13B. The size of LLaVA in Equation (2) is 13B and ChatGPT in Equation (3) is gpt-3.5-turbo-1106.
To verify the effectiveness of our proposed FGAIF, we compare our method with baselines on several benchmarks, including QA-based hallucination benchmarks POPE (Li et al., 2023d) and MMHal-Bench (Sun et al., 2023), hallucination metrics CHAIR (Rohrbach et al., 2018) and FaithScore (Jing et al., 2024), and the general benchmark LLaVA-Bench (Liu et al., 2023c).
CHAIR is a framework to quantify object hallucination in image captions. This method compares objects generated in captions against the ground truth objects within the images. CHAIR assesses hallucination on two levels: sentence-level and instance-level. The sentence-level score, referred to as CHAIRS, quantifies the proportion of captions that contain hallucinated content, whereas the instance-level score, CHAIRI, measures the frequency of hallucinated objects relative to the total number of objects mentioned by the model. Our evaluation involves a randomly selected subset of 1,000 images from the MSCOCO validation set, allowing for an analysis of our model’s performance in minimizing object existence hallucination.
FaithScore is another framework to assess the accuracy and relevance of response generated by LVLMs. This innovative approach focuses on evaluating the consistency of atomic facts within the response against the depicted scenes in the input images. Different from CHAIR, FaithScore can demonstrate the model’s hallucination performance in terms of object existence, attribute, and relation. Our evaluation involves a randomly selected subset of 1,000 images from the MSCOCO validation set, allowing for an analysis of our model’s performance in mitigating object existence, attribute, and relation hallucination. It also provides an instance-level score F-Score and sentence-level score F-ScoreS.
@@ 5.1 Experimental Details
CHAIRI | CHAIRS | F-Score | F-ScoreS | ||
MiniGPT-47B | 9.4 | 17.4 | 63.9 | 61.8 | 245.1 |
mPLUG-Owl7B | 6.2 | 9.5 | 85.6 | 65.7 | 75.2 |
InstructBLIP7B | 2.4 | 3.8 | 93.6 | 80.0 | 45.6 |
InstructBLIP13B | 2.7 | 4.0 | 94.1 | 80.8 | 46.3 |
LLaVA7B | 9.1 | 22.0 | 88.9 | 72.3 | 216.0 |
LLaVA13B | 10.3 | 19.8 | 87.9 | 68.3 | 121.0 |
LLaVA-RLHF7B | 4.6 | 7.0 | 89.3 | 71.1 | 58.8 |
LLaVA-RLHF13B | 7.7 | 20.3 | 89.7 | 73.8 | 413.8 |
FGAIF7B | 3.9 | 6.2 | 91.2 | 74.7 | 60.2 |
@@
CHAIRI | CHAIRS | F-Score | F-ScoreS | F1 | Over | Hal | |
FGAIF7B | 3.9 | 6.2 | 91.2 | 74.7 | 83.4 | 3.09 | 0.36 |
@@
5.2 On Model Comparison (RQ1)
The results on QA-based hallucination benchmarks (i.e., POPE and MMHal-Bench) are summarized in Table 1 and Table 2. From this table, we have several observations. (1) LLaVA7B and InstructBLIP7B performs worse than LLaVA13B and InstructBLIP13B on most cases, respectively. Compared with LLaVA13B, LLaVA7B has a strong hallucination problem, especially its over-confident problem on POPE. This indicates the importance of model size. (2) LLaVA-RLHF7B is better than LLaVA7B, which indicates the superiority of further fine-tuning with human feedback. Notably, LLaVA-RLHF7B even has a better performance compared to LLaVA13B, even though the latter has specifically more parameters. (3) Our model consistently performs better than the previous advanced in terms of most metrics and testing sets. This verifies that fine-grained artificial intelligence feedback also can be beneficial for hallucination mitigation in LVLMs. (4) Our FGAIF surpasses LLaVA-RLHF7B across all metrics. This implies the advantage of fine-grained artificial intelligence feedback compared to human feedback. (5) To further understand the performance of our FGAIF, we split the MMHal-Bench into three classes based on the original dataset: a) object existence (class “adversarial object”), b) object attribute (classes “object attribute” and “counting”), and c) object relation (class “spatial relation”). We observe that our method consistently achieves the best performance across all question categories.
We further show the performance of our FGAIF and baselines on hallucination metrics CHAIR and FaithScore in Table 3. InstructBLIP7B and InstructBLIP13B achieve the best performance in CHAIR and FaithScore metrics. The potential reason is that these two models tend to generate short answers and these two metrics just measure the precision of faithfulness but do not contain recall of faithfulness. Despite this, our FGAIF still outperforms the RLHF-based baseline (i.e., LLaVA-RLHF7B) whose answers are shorter than FGAIF, which verifies the superiority of our method.
In addition, Table 2 shows the comprehensive performance comparison of our FGAIF and the baseline methods on the general benchmark LLaVA-Bench. From this table, we observed that most models perform worst on the “Detail” (i.e., detailed description) subset and perform best on the “Complex” (i.e., complex questions) subset. This may be due to the reason that the “Detail” (i.e., detailed description) subset has more stringent requirements for faithfulness because all the content of the response is required to be an accurate description of the input image. On the contrary, the “Complex” (i.e., complex questions) subset often explores the extended content of an image, sometimes leading to open-ended discussions. Therefore, the demand for strict consistency with the image isn’t as critical. In addition, we found that the RLHF can boost the LVLM’s performance on the whole LLaVA-Bench from 81.0 (LLaVA7B) to 94.1 (LLaVA-RLHF7B). Furthermore, our FGAIF can bring more performance gain in terms of the “Conv” subset, “Detail”, “Complex” subset, and full set), compared with LLaVA-RLHF7B. This further indicates the advance of our method.
5.3 On Ablation Study (RQ2)
To verify the effect of each component in our FGAIF, we devise the following variant methods for ablation study:
w/o-Obj: To demonstrate the effect of the object hallucination feedback, we remove the object existence reward model in this method;
w/o-Att: To show the necessity of the attribute hallucination feedback, we remove the object attribute reward model in this method;
w/o-Rel: To demonstrate the effect of the relation hallucination feedback, we remove the object relation reward model in this method;
w/o-AIF: To show the benefit of using reinforcement learning from fine-grained artificial intelligence feedback, we remove all the reinforcement learning components in this variant;
w-Coarse: To verify the advance of the fine-grained feedback compared with the traditional coarse-grained uni reward model, we replace the three fine-grained reward models with one reward model which also is trained with AI annotated data and the training phase is the same as the previous work (Sun et al., 2023).